53 5 1 2 A. 6x-Sy= 16 C. D. Mark this and return
Answer #1

We will find this by Slope intercept form.

If you know the slope m , and y -intercept (0,b) of a line (the point where the line crosses the y -axis), you can write the equation of the line in slope-intercept form .

      y=mx+b....................(1)

where m is the slope.

To find the slope, use the two points (x1,y1) and (x2,y2) and substitute these in the slope formula:

m=(y2?y1 )/(x2?x1)

In our case the y-intercept (0,-2) of line is -2 i.e. b= -2

Now we see from figure that (4,1) and (0,-2) are two points on the line.

so slope m = (-2-1)/(0-4)

= -3/-4

= 3/4

Putting these values in slope intercept form i.e in equation (1)

So we have

y= (3/4)x -2

So option A is correct.